OpenStudy (anonymous):

who is doing Geometry Part I, Section 2 part 2 for high school for ashworth college/james madison high school then can you help me with them all I really need help.

OpenStudy (anonymous):

an part 2

OpenStudy (anonymous):

can anyone that has done the Geometry part one and two could you please help me out

OpenStudy (wolf1728):

Maybe if you just stated a geometry problem you'd get a lot more responses than just students at James Madison High School. Anyway, please post the problem.

OpenStudy (anonymous):

Identify the hypothesis and conclusion of this conditional statement: If two lines intersect at right angles, then the two lines are perpendicular. A. Hypothesis: The two lines are perpendicular. Conclusion: Two lines intersect at right angles. B. Hypothesis: Two lines intersect at right angles. Conclusion: The two lines are perpendicular. C. Hypothesis: The two lines are not perpendicular. Conclusion: Two lines intersect at right angles. D. Hypothesis: Two lines intersect at right angles. Conclusion: The two lines are not perpendicular. What is a counterexample for the conjecture? Conjecture: The product of two positive numbers is greater than the sum of the two numbers. A. 3 and 5 B. 2 and 2 C. A counterexample exists, but it is not shown above. D. There is no counterexample. The conjecture is true. Given: X/5 + 9 =11 Prove: x=10 x/5 + 9 = 11 a.___________ x/5 = 2 b.____________ x=10 c.____________ A. a. Given b. Subtraction Property of Equality c. Division Property of Equality B. a. Given b. Subtraction Property of Equality c. Multiplication Property of Equality C. a. Given b. Addition Property of Equality c. Multiplication Property of Equality D. a. Given b. Addition Property of Equality c. Division Property of Equality What is a counterexample for the conjecture? Conjecture: Any number that is divisible by 4 is also divisible by 8. A. 24 B. 40 C. 12 D. 26 Which biconditional is NOT a good definition? A. A whole number is even if and only if it is divisible by 2. B. A whole number is odd if and only if the number is not divisible by 2. C. An angle is straight if and only if its measure is 180. D. A ray is a bisector of an angle if and only if it splits the angle into two angles. theses are the ones I need help with

OpenStudy (wolf1728):

Wow that's a LOT of questions. I'll answer the first one. (Wish they were numbered instead of everything being lettered A, B, C, D) The first question about right angles A & B are true.